One of two methods must be used to produce expansion anchors. Method A costs $50,000 initially and will have a $20,000 salvage value after 3 years. The operating cost with this method will be $29,000 per year. Method B will have a first cost of $105,000, an operating cost of $20,000 per year, and a $38,000 salvage value after its 3-year life. The interest rate for both the methods is 15%.

Which method should be used on the basis of a present worth analysis?

The present worth of method A is $  and that of method B is $  .
Answer #1

Present worth of A = -50000 - 29000(P/A, 15%, 3) + 20000(P/F, 15%, 3)

= -50000 - 29000*2.28323 + 20000*0.65752

= -103,063

Present worth of B = -105000 - 20000(P/A, 15%, 3) + 38000(P/F, 15%, 3)

= -105000 - 20000*2.28323 + 38000*0.65752

= -125,679

Since the PW of A is low (it is less expensive), we select A.